# cadrille: Multi-modal CAD Reconstruction with Online Reinforcement Learning

### v2

@@ Appendix C Implementation Details
Our model is built on top of Qwen2-VL wang2024qwen2-vl and uses its native capabilities of image and text understanding. In all experiments on multi-view image CAD reconstruction, we use four images. Images are rendered with fixed camera positions, and concatenated into 2x2 grid, forming a combined image of size 268 $\times$ 268 px (as shown in Fig. 7 and  9). This combined image is passed through the Qwen vision encoder, that outputs 400 input tokens. The point cloud injecting into LLM is implemented exactly as in CAD-Recode rukhovich2024cad-recode . Specifically, our input consists of 256 unordered 3D points without normals, sampled from the surface using the furthest point sampling method. The points are projected into shared embedding space with a single linear layer.
@@
The RL fine-tuning hyperparameters are listed in Tab. 10 (DPO) and  11 (Dr. CPPO). In each experiment, the model is initialized with the weights of an SFT model trained on point clouds and images, and then fine-tuned only on images. All fine-tuning experiments are performed on 8 H100 GPUs.
@@ Appendix D Miscellaneous
CC3D scans
Following CAD-Recode, we treat experiments with CC3D dataset as an emulation of a real-world experiment. CC3D contains scans of the physical objects paired with ground truth CAD models. As shown in Fig. 11, CC3D scans contain artifacts such as surface noise, smoothed edges, and missing parts. In our experiments, we sample points from the surfaces of these real noisy scans instead of the perfect CAD models, which essentially makes the experimental scenario much more realistic.
Besides, CC3D contains generally more complex CAD models, which are constructed using operations beyond simple extrusion, including revolution, chamfer, and fillet.

### v3

@@ Appendix D Implementation Details
Our model is built on top of Qwen2-VL Wang et al. (2024) and uses its native capabilities of image and text understanding. In all experiments on multi-view image CAD reconstruction, we use four images. Images are rendered with fixed camera positions, and concatenated into 2x2 grid, forming a combined image of size 268 $\times$ 268 px (as shown in Fig. 8 and  9). This combined image is passed through the Qwen vision encoder, that outputs 400 input tokens. The point cloud injecting into LLM is implemented exactly as in CAD-Recode. Specifically, our input consists of 256 unordered 3D points without normals, sampled from the surface using the furthest point sampling method. The points are projected into shared embedding space with a single linear layer.
@@
The RL fine-tuning hyperparameters are listed in Tab. 13 (DPO) and  14 (Dr. CPPO). In each experiment, the model is initialized with the weights of an SFT model trained on point clouds and images, and then fine-tuned only on images. All fine-tuning experiments are performed on 8 H100 GPUs.
@@ Appendix E Miscellaneous
CC3D scans
Following CAD-Recode, we treat experiments with CC3D dataset as an emulation of a real-world experiment. CC3D contains scans of the physical objects paired with ground truth CAD models. As shown in Fig. 11, CC3D scans contain artifacts such as surface noise, smoothed edges, and missing parts. In our experiments, we sample points from the surfaces of these real noisy scans instead of the perfect CAD models, which essentially makes the experimental scenario much more realistic. Besides, CC3D contains generally more complex CAD models, which are constructed using operations beyond simple extrusion, including revolution, chamfer, and fillet.